OpenStudy (anonymous):

2 x (3 x 5) Rewrite the expression below to demonstrate the associative property of multiplication.

OpenStudy (anonymous):

\[(2\times 3)\times 5\]

OpenStudy (anonymous):

2*(3*5) = 2*15 = 30 OR (2*3)*5 = 6*5 = 30 In whatever way you multiply the numbers, you get 30. So the expression demonstrates the associative property of multiplication.

OpenStudy (anonymous):

you can't "demostrate" that property, the associative property is an axiom of the real number with the multiplication. axiom is a property, or sentence, accepted as true, without demostration.